OpenStudy (anonymous):

what are 2 equivalent fractions of 2/3

OpenStudy (anonymous):

3/4 and 4/5

OpenStudy (anonymous):

To get the answer, you can multiply 2 and 3 by any number and the resulting ratio will be the equivalent.

OpenStudy (anonymous):

3/4 and 4/5 are WRONG!

OpenStudy (anonymous):

Let us multiply 2/3 by 2/2. So, you get 4/6. It is same as 2/3. Let us multiply 2/3 by 5/5. So, you get 10/15. It is same as 2/3.

OpenStudy (anonymous):

ok im sorry by the way

OpenStudy (jamesj):

Another example, slightly more exotic but important is: \[ \frac{2}{3} = \frac{2}{3} \times \frac{-1}{-1} = \frac{2 \times -1}{3 \times -1} = \frac{-2}{-3} \]